OpenStudy (anonymous):

Can anyone help me with Circle in the Coordinate Plane?......i cant wrap my head around this stuff to save my life ive been going over the book and quiz i need for days but nothing tells me how to solve the questions.

terenzreignz (terenzreignz):

Sure... can you be more specific? :)

terenzreignz (terenzreignz):

If you could post a specific question, for an example, I'll see what I can do~

OpenStudy (anonymous):

The center of a circle is (h,7) and the radius is 10. the circle passes through(3,-1). Find all possible values of h.

terenzreignz (terenzreignz):

Okay... you know how to get the equation of a circle, when you're given its center (h,k) and its radius r?

OpenStudy (anonymous):

ummmmm.....i may but i wanna be sure....this is important

terenzreignz (terenzreignz):

Okay :) If you're given the center \(\large (\color{red}h \ , \ \color{green}k)\) of a circle as well as its radius \(\large \color{blue}r\) then the equation of the circle is \[\Large (x-\color{red}h)^2 + (y-\color{green}k)^2=\color{blue}r^2\]

OpenStudy (anonymous):

ok......so how should the plug ins be placed?

terenzreignz (terenzreignz):

Well, what are your values for h, k, and r?

OpenStudy (anonymous):

h is what you need. k is 7 and r is 10?

terenzreignz (terenzreignz):

That's correct :D \[\Large (x-\color{red}h)^2 + (y-\color{green}7)^2=\color{blue}{10}^2\] \[\Large (x-h)^2 + (y-7)^2 = 100\] Catching me so far?

OpenStudy (anonymous):

ok.....now how does the (3,-1) fit in? im guessing thats the x,y?

terenzreignz (terenzreignz):

YES Exactly :D \[\Large (\color{red}3-h)^2+(\color{blue}{-1}-7)^2=100\] Are you sure you needed help with this? :3

OpenStudy (anonymous):

well ive been having problems with geo all year and this counts for alot since i got my final dropped so need to be absolutly sure i know what im doing

terenzreignz (terenzreignz):

Okay, so can you solve that ^ equation now? It's just a quadratic equation in one variable... use whatever method suits you... but if you get stuck, call me ;)

OpenStudy (anonymous):

i certainly will thank you so very much

terenzreignz (terenzreignz):

No problem :)

OpenStudy (anonymous):

terenzreignz .....what determins the symbol inside the perenthesis?

terenzreignz (terenzreignz):

what?

OpenStudy (anonymous):

the symbol inside ( )....like - or +

terenzreignz (terenzreignz):

I'm not getting, need to be more specific...

OpenStudy (anonymous):

the standerd equation is (x-h)^2+(y-k)=r^2....what would make be (x+h)^2+(y-k=r^2 or the other way around?

terenzreignz (terenzreignz):

not x+h but x-h

OpenStudy (anonymous):

yeah what would determine the difference

terenzreignz (terenzreignz):

what do you mean by "the other way around" ?

OpenStudy (anonymous):

either (x-h)+(y-k)....(x+h)+(y-k)....or (x-h)+(y+k)

terenzreignz (terenzreignz):

Oh... if the center has positive coordinates, then (x-h) and (y-k) if the center has negative coordinates, then you'd have pluses... for instance, if the center was (-4 , 2) then it'd be (x+4) and (y-2)

OpenStudy (anonymous):

ohhhhhhhhhhhhhhhhhhhhh....ok thank you much.....makes sense now

OpenStudy (anonymous):

sorry i confused you

terenzreignz (terenzreignz):

:)